# IBIS: A Powerful Hybrid Architecture for Human Activity Recognition

### v1

@@ Abstract
The increasing interest in Wi-Fi sensing stems from its potential to capture environmental data in a low-cost, non-intrusive way, making it ideal for applications like healthcare, space occupancy analysis, and gesture-based IoT control. However, a major limitation in this field is the common problem of overfitting, where models perform well on training data but fail to generalize to new data. To overcome this, we introduce a novel hybrid architecture that integrates Inception-BiLSTM with a Support Vector Machine (SVM), which we refer to as IBIS. Our IBIS approach is uniquely engineered to improve model generalization and create more robust classification boundaries. By applying this method to Doppler-derived data, we achieve a movement recognition accuracy of nearly 99%. Comprehensive performance metrics and confusion matrices confirm the significant effectiveness of our proposed solution.
I Introduction
@@
To address these limitations, we propose a novel hybrid neural network architecture named Inception-BiLSTM (IBIS). The Inception component efficiently extracts local features from the input data [5], while the Bidirectional Long Short-Term Memory (BiLSTM) component captures long-range temporal dependencies. To further enhance the model’s predictive power and refine classification boundaries, we integrate a SVM with a non-linear kernel in a post-processing stage. This ensemble learning approach [6] leverages the strengths of multiple algorithms to achieve a more robust and reliable solution.
Our model’s effectiveness was evaluated on five and eight human activity classification tasks using Doppler-derived data. We compared our IBIS approach against two state-of-the-art methods: the SHARP algorithm [7], a recent ensemble-based method for HAR, and an Attention-Based BiLSTM (ABLSTM) approach [8] that uses raw CSI. The results demonstrate that our hybrid Inception-BiLSTM network, augmented with a non-linear SVM, significantly outperforms both benchmark models. Our method achieved an average accuracy of nearly 99%, validating the benefits of combining Doppler features with a sophisticated hybrid deep learning architecture for enhanced HAR performance, particularly in challenging generalization scenarios.
The remainder of this paper is structured as follows. Section II presents a comprehensive overview of existing research on HAR using hybrid neural networks, providing a fresh perspective on the field. Section III details the development and implementation of our proposed algorithms, including the specific neural network architecture. The validation process, which includes activity classification confusion matrices and Receiver Operating Characteristic (ROC) curve analysis, is described in Section IV. Finally, Section V provides a summary of our findings and offers concluding remarks.
II Related Works
This section provides a review of state-of-the-art research in Wi-Fi sensing, focusing on the machine learning algorithms commonly used for CSI data analysis. We also discuss ensemble learning techniques that have been applied to optimize model performance. Recent studies have explored various approaches to enhance Wi-Fi sensing capabilities. Table I presents a comparative overview of the models and key features from previous studies on CSI-based Human Activity Recognition.
Son et al. [9] improved indoor occupancy detection by proposing a preprocessing method that uses nonlinear transformations to expand CSI data into three channel vectors, which amplified signal differences. A CNN was then applied, achieving an average accuracy of approximately 95%. This work highlights the importance of effective signal preprocessing for boosting classification performance.
Ensemble learning has emerged as a key strategy to improve the reliability and accuracy of machine learning models. Wei et al. [10] used an ensemble learning approach for HAR by combining feature extraction with an SVM classifier. This method reportedly outperformed other literature approaches, reaching an accuracy of about 98.95%.
@@ II Related Works
The SHARP algorithm, as proposed by Meneghello et al. [7], provides a foundational method for HAR based on Doppler traces. Their work addressed noise and hardware interference and, after applying a neural network-based refinement, achieved an accuracy of up to 95%. The authors also demonstrated that antenna merging could further enhance accuracy. Building upon this, Cominelli et al. [12] investigated the impact of Wi-Fi 6 features and developed a large-scale dataset for benchmarking. However, they reported difficulties in reproducing the SHARP algorithm’s results, suggesting a sensitivity to environmental variability. Their findings also indicated that SHARP’s performance degrades with increased bandwidth and a higher number of activity classes.
Despite these advancements, a major challenge in Wi-Fi sensing remains the issue of poor generalization to unseen data. He et al. [8] addressed this limitation by introducing a cross-subject transfer learning method based on a CNN-ABLSTM model. By training on separate source and target domains, they achieved an accuracy of 85.82% in cross-subject recognition tasks, demonstrating a promising path for creating more robust and generalizable models.
TABLE I: Comparison of Models and Features in Human Activity Recognition Using CSI
Specification [9] [10] [7] [12] [11] [8]
Model CNN CNN+RF+SVM Inception Inception DNN-ABLSTM CNN-ABLSTM
Accuracy 95% 98.85% 88.12% 77.72% 92.06% 85.82%
Application Indoor Occupancy HAR HAR HAR HAR HAR
Network 802.11 802.11n 802.11ac 802.11ac/ax 802.11n -
Bandwidth 20MHz - 80MHz 80/160MHz - -
CSI Data Amplitude Temporal and Spatial Correlation Phase CFR Phase CFR Amplitude and Phase Raw CSI
III System Modeling
III-A Dataset
@@
The dataset for this research is based on the SHARP study [7] and consists of CSI data collected from three volunteers (one male, two females) performing eight activities in various indoor environments. These activities: empty, sitting, stand up, walking, running, jumping, and arm gym—were each recorded for 120 seconds. The data was gathered in a bedroom, living room, and university laboratory during two periods: April–December 2020 and January 2022. A Nexmon CSI Extraction Tool and an Asus RT-AC86U Wi-Fi router were used to extract the CSI from Wi-Fi 2.4 GHz preamble headers.
@@
III-B Decision-Making Process and Algorithm Steps
@@
To understand the purpose of our work, it is essential to first comprehend the original SHARP method. The SHARP process begins with the collection of CSI data from three individuals, with each of the eight activities sampled for 120 seconds. After the CSI data is extracted from the Wi-Fi 2.4 GHz packet preambles, it undergoes three key stages, as shown in Figure  1:
  1. 1.
@@
    Pre-processing: This stage sanitizes the CSI data by eliminating noise, interference, and phase ambiguity. A reference signal path is used to remove phase offsets, recovering a clean signal that reveals the Doppler traces associated with movement;
  2. 2.
@@
    Classification: The sanitized Doppler traces are fed into an Inception neural network. To improve generalization beyond the original architecture’s spatial feature learning, the network was enhanced with a BiLSTM layer, creating a hybrid Inception-BiLSTM model. This modification allows the model to capture both spatial and temporal dependencies, significantly boosting performance in human activity recognition;
  3. 3.
    Post-processing: The original SHARP method used a majority voting mechanism among the four receiving antennas to determine the final activity label. This step smoothed outliers and improved robustness.
III-C Hybrid Inception-BiLSTM (IBIS) Network
To interpret and classify data from Doppler traces, we designed a hybrid neural architecture combining two networks. The Inception network processes global features and learns spatial characteristics, while the BiLSTM network captures temporal dependencies and patterns. This combination is designed to optimize both classification performance and generalization ability.
@@ III-C Hybrid Inception-BiLSTM (IBIS) Network
Figure 2 illustrates the integrated architecture of these two neural structures. While some layers like Normalization, Reshape, and Dropout are omitted for visualization, they are described in the text. Table II provides the input and output dimensions and parameter counts for each block. The shaded and hatched purple areas in the figure highlight the novel components of our proposal. Our approach enhances the standard Inception network, which typically only includes activation and convolution layers for learning local features. We added a MaxPooling layer and Batch Normalization to improve convergence and filter relevant features. Furthermore, a BiLSTM network was incorporated to enable the model to learn temporal patterns and motion dynamics from Doppler shifts.
III-C1 Inception Network
@@
The Inception network is configured to extract spatial features from the Doppler trace data. A description of each block and its function is as follows:
@@
III-C2 BiLSTM Network
The second part of the architecture is the BiLSTM network, which was included to capture crucial temporal patterns and dependencies from the Doppler traces. These traces are time-series signals that reflect movement and channel variations, often exhibiting complex patterns that the network must recognize.
@@ III-C2 BiLSTM Network
  • Dense Layer: A fully connected layer that generates the final output, with the number of units depending on the number of output classes (Table II).
While the CNN-ABLSTM model from a comparative study [8] has a lower computational cost due to its lighter architecture, it only achieves an accuracy of 85.82%, compared to the 98.95% of our IBIS method. This performance gap is attributed to the limited number of attention layers in their model, which are insufficient to capture the complex temporal and spatial patterns and subtle variations present in data with high variability and Doppler effects. Our hybrid Inception-BiLSTM network, with its greater depth and robust attention mechanisms, is better equipped to handle complex local features, leading to higher learning capacity and a more efficient convergence rate. This is because the Inception module learns spatial features, while the BiLSTM module captures temporal dependencies, making the combined model more robust to noise and interference, which is ideal for classifying Doppler traces.
Figure 2: Architecture of the IBIS hybrid neural network for Human Activity Recognition using Doppler traces, showing all convolutional and attention layers.
TABLE II: Inception-BiLSTM Architecture (Compact Notation)
@@
III-D Support Vector Machine with Radial Basis Function Kernel
The hybrid IBIS network is effective at detecting and classifying both global patterns and temporal sequences. However, for accurate classification of CSI-derived data, this architecture alone is insufficient. To address this, a SVM was integrated into the post-processing stage as an ensemble learning strategy. Unlike common approaches where SVM is the primary classifier, this work uses it to refine the final classification results and improve generalization. The SVM’s ability to handle outliers and its use of decision margins contribute to better model generalization, which is particularly useful for Doppler trace data that is prone to noise and distortions. This allows the SVM to refine the final classification results, especially for samples that might be misclassified by the neural network.
A key advantage of SVM is its kernel function flexibility. The Radial Basis Function (RBF) kernel [13] [14] was chosen for this study’s Doppler data due to its ability to adapt decision boundaries to non-linear distributions. The penalty parameter C was set to 1, providing an optimal balance between training accuracy and generalization. Additionally, class balancing was applied to the SVM configuration to ensure that each class was given equal weight, which prevents the model from being biased toward the majority class and leads to a fairer, more robust classification. Figure 3 illustrates the SVM model’s application, showing five distinct classes separated by well-defined boundaries. Proper class segmentation is essential for understanding the unique features of each class, leading to improved classification accuracy.

### v2

# IBIS: A Hybrid Inception-BiLSTM and SVM Ensemble for Robust Doppler-based Human Activity Recognition

@@ Abstract
Wi-Fi sensing is a leading technology for Human Activity Recognition (HAR), offering a non-intrusive and cost-effective solution for healthcare and smart environments. Despite its potential, existing methods struggle with domain shift issues, often failing to generalize to unseen environments due to overfitting. This paper proposes IBIS, a robust ensemble framework combining Inception-Bidirectional Long Short-Term Memory (BiLSTM) for feature extraction and Support Vector Machine (SVM) for classification of Doppler signatures. The proposed architecture specifically targets generalization capabilities. Experimental results on multiple datasets show that IBIS achieves 95.40% accuracy, delivering a 7.58% performance gain compared to standard architectures in cross-scenario evaluations on external datasets. The analysis confirms that IBIS effectively mitigates environmental dependency in Wi-Fi-based HAR.
{IEEEkeywords} IBIS, Channel State Information, Human Activity Recognition, Support Vector Machine, Inception, BiLSTM, Doppler traces, Ensemble Learning.
1 Introduction
@@
Recent studies have explored various approaches to enhance Wi-Fi sensing capabilities. Son et al. [5] improved indoor occupancy detection by proposing a pre-processing method that uses nonlinear transformations to expand CSI data into three channel vectors, which amplified signal differences. A CNN was then applied, achieving an average accuracy of approximately 95%. This work highlights the importance of effective signal pre-processing for boosting classification performance.
Ensemble learning has emerged as a key strategy to improve the reliability and accuracy of machine learning models. Wei et al. [6] used an ensemble learning approach for HAR by combining feature extraction with an SVM classifier.
@@
The SHARP algorithm, as proposed by Meneghello et al. [8], provides a foundational method for HAR based on Doppler traces. Their work addressed noise and hardware interference and, after applying a neural network-based refinement, achieved an accuracy of up to 95%. The authors also demonstrated that antenna merging could further enhance accuracy. Building upon this, Cominelli et al. [9] investigated the impact of Wi-Fi 6 features and developed a large-scale dataset for benchmarking. However, they reported difficulties in reproducing the SHARP algorithm’s results, suggesting a sensitivity to environmental variability. Their findings also indicated that SHARP’s performance degrades with decreased bandwidth and a higher number of activity classes.
Despite recent advances, Wi-Fi sensing still faces limited generalization to unseen data. Prior work, such as the CNN-ABLSTM cross-subject transfer learning method by He et al. [10], achieved 85.82% accuracy but highlighted the need for more robust models. Moreover, most existing studies rely only on the CSI amplitude, overlooking phase information, which is highly sensitive to noise and interference.
To address these gaps, we propose IBIS, a hybrid Inception-BiLSTM architecture combined with a non-linear SVM in the post-processing stage. Unlike conventional ensemble approaches that fuse models during training, IBIS leverages the probability outputs of internal neural layers as high-level feature vectors for the SVM, enabling sharper decision boundaries. Our framework operates solely on Doppler traces extracted from the CSI phase component, exploiting information typically ignored in the literature.
Expanding the introductory analysis presented in [11], IBIS was evaluated on five- and eight-class activity recognition tasks and compared against two state-of-the-art baselines: the SHARP algorithm [8] and an ABLSTM model using raw CSI [10]. To assess robustness across environments, we also included a complementary Doppler-based dataset [9]. The proposed method achieved an average accuracy of 95.40%, significantly outperforming all benchmark models and demonstrating strong generalization capabilities across different datasets and noise conditions.
The main contribution of this proposal is to improve the generalization capability of CSI data with respect to HAR by implementing a combination of spatial features and temporal sequences through the application of the proposed framework. This approach enables higher activity recognition accuracy. In addition, the inclusion of a SVM algorithm in the post-processing stage further enhances the framework, thereby demonstrating its effectiveness compared to traditional deep learning–based methods.
Table 1: Survey of Models and Features in Using CSI
Method / Authors Year Task Model Bandwidth Network Accuracy CSI Data
Son et al. [5] 2024 Indoor Ocupancy CNN 20 MHz 802.11 95% Amplitude
WiAres [6] 2020 HAR CNN+RF+SVM - 802.11n 98.85% Temporal / Spatial Correlation
SHARP [8] 2023 HAR Inception 20/40/80 MHz 802.11 88.12% Phase CFR
Cominelli et al. [9] 2023 HAR Inception 20/40/80/160 MHz 802.11ac/ax 77.12% Phase CFR
Ye et al. [7] 2024 HAR DNN-ABLSTM - 802.11n 92.06% Amplitude and Phase
He et al. [10] 2025 HAR CNN-ABLSTM - - 85.82% Raw CSI
Widar [12] 2022 Gesture Recognition CNN+RNN - - 92.7% Amplitude
EfficientFi [13] 2022 HAR CNN 40 MHz 802.11ax 98% Amplitude
WiGr [14] 2022 Gesture Recognition CNN+LSTM 40 MHz 802.11ax 86.8%–92.7% Amplitude
Wi-SensiNet [15] 2024 HAR CNN+BiLSTM 20 MHz 802.11 99% Amplitude
IBIS* 2025 HAR Inception-BiLSTM+SVM 80/160 MHz 802.11ac/ax 95.40% Phase CFR
Note: The symbol * represents our proposal Inception - BILSTM + SVM architecture
It is important to highlight that our proposal achieved better results compared to the related works presented in Table 1. In terms of accuracy, the IBIS framework outperformed the works that use only amplitude. The main point is that when using phase information, it must be carefully handled to avoid possible classification errors caused by interference and existing offsets.
The remainder of this paper is organized as follows. Section 2 presents the definition of CSI as well as a detailed description of the datasets used in this study. Section 3 describes the development and implementation of the proposed algorithms, including the specific neural network architecture. The validation process, encompassing activity classification confusion matrices and Receiver Operating Characteristic (ROC) curve analysis, is presented in Section 4. Finally, Section 5 summarizes the main findings and provides concluding remarks.
2 Channel State Information Data Description
This section presents the definition of CSI as adopted in IEEE 802.11 standard networks and also describes the datasets used in this proposal for data collection.
2.1 Channel State Information
CSI is a data collection system that describes how a wireless signal propagates from the transmitter to the receiver. CSI captures the signal propagation characteristics in physical environments, including the combined effects of diffraction, reflection, and scattering [5].
Adopting the IEEE 802.11 standard, it employs advanced techniques such as Multiple-Input Multiple-Output (MIMO) and Orthogonal Frequency Division Multiplexing (OFDM) at the physical layer. These technologies aim to improve the supported data capacity and enhance channel orthogonality in environments affected by multipath propagation, which occurs when transmitted signals between the transmitter and receiver follow different paths due to phenomena like collisions and external interference.
For each transmitter/receiver antenna pair, CSI provides detailed information about the phase and amplitude attenuation experienced by the signal through multiple paths for each subcarrier. Specifically, the CSI for each subcarrier $k$ between a transmitter with $N_{t}$ antennas and a receiver with $N_{r}$ antennas is represented by the matrix $\mathbf{H}(k)\in\mathbb{C}$ [5]:
$\mathbf{H}(k)=\begin{bmatrix}h_{11}(k)&\cdots&h_{1N_{t}}(k)\\ h_{21}(k)&\cdots&h_{2N_{t}}(k)\\ \vdots&\ddots&\vdots\\ h_{N_{r}1}(k)&\cdots&h_{N_{r}N_{t}}(k),\end{bmatrix},k\in[1,K]$ (1)
where $K$ represents the number of subcarriers used, and $h_{rc}(k)$ denotes the complex channel gain from the $c{\text{-th}}$ transmitter antenna to the $r{\text{-th}}$ receiver antenna. Each element $h_{rc}(K)$ in the CSI matrix $\mathbf{H}(k)$ can be decomposed into amplitude and phase components, as described by the following:
$h_{rc}(k)=|h_{rc}(k)|\cdot\exp\left(j\theta_{rc}(k)\right),$ (2)
where $\left|h_{rc}(k)\right|$ is defined as the amplitude component, and $\theta_{rc}(k)$ is the phase component of $h_{rc}(k)$.
Since CSI contains both amplitude and phase information for each subcarrier and for each pair of transmit and receive antennas, it provides high-resolution data compared to simple measurements such as RSS (Received Signal Strength). In this way, CSI data can be considered as a form of ”imaging” of the signal propagation environment.
Changes in the environment, such as the presence or movement of people, directly affect Wi-Fi signal propagation. These alterations in the signal path also lead to variations in CSI. Since different people create distinct patterns, it is possible to estimate the number of individuals present by learning these patterns—this is the key concept behind people counting systems based on CSI-extracted data [5].
2.2 Dataset A
@@
One of the datasets used for this research is based on the SHARP study [8] and consists of CSI data collected from three volunteers (one male, two females) performing eight activities in various indoor environments. These activities: empty, sitting, stand up, walking, running, jumping, and arm gym, were each recorded for 120 seconds. The data was gathered in a bedroom, living room, and university laboratory during two periods: April–December 2020 and January 2022. A Nexmon CSI Extraction Tool and an Asus RT-AC86U Wi-Fi router were used to extract the CSI from Wi-Fi 2.4 GHz preamble headers.
@@
2.3 Dataset B
For application in the experiment, we considered using a second dataset, developed by Cominelli et al. [9]. Although originally designed for IEEE 802.11ax (Wi-Fi 6), thereby expanding the application scope. The main characteristics of this dataset are summarized below.
This dataset was developed using four Asus RT-AX86U routers interconnected to form a local-area network (LAN) and controlled by a host notebook. One of the routers is configured to generate dummy Wi-Fi traffic at a constant rate through feature injection for AX-CSI, while the other three extract the corresponding CSI. According to the authors, due to the use of IEEE 802.11ax, data grows very rapidly, making it necessary to store the recordings in PCAP files.
Data collection was performed across seven different scenarios:
  • S1 to S5: Conducted in the same physical environment, where neither the receivers nor the transmitter were moved during the entire duration of the experiments;
  • S6: An office environment containing chairs and desks;
  • S7: A large hall, representing an open environment without obstacles.
Twelve distinct activities were recorded: Walk, Run, Jump, Sitting, Empty room, Standing, Wave hands, Clapping, Lay down, Wiping, Squat, Stretching with a sampling rate of 80 samples per second.
For our experiment, scenario S2 was selected. This environment consists of a laboratory with reflective surfaces, which makes prediction more difficult.
Due to the accuracy challenges encountered with the original model, our intention is to apply the IBIS framework in different environments and scenarios in order to evaluate its behavior and robustness.
3 System Modeling
In this section, the main definitions related to the IBIS framework are presented, including the architectural composition, as well as the derived pseudocode outlining the main steps of the framework operation.
3.1 Decision-Making Process and Algorithm Steps
@@
To understand the purpose of our work, it is essential to first comprehend the SHARP method. After the CSI data is extracted from the Wi-Fi 802.11 ac/ax packet preambles, it undergoes three key stages, as shown in Figure  1:
3.1.1 Pre-processing
@@
This stage sanitizes the CSI data by eliminating noise, interference, and phase ambiguity. A reference signal path is used to remove phase offsets, recovering a clean signal that reveals the Doppler traces associated with movement.
3.1.2 Classification
@@
The sanitized Doppler traces are fed into an Inception neural network. To improve generalization beyond the original architecture’s spatial feature learning capabilities, the network was enhanced with a BiLSTM layer, creating a hybrid Inception–BiLSTM model. This modification enables the architecture to capture both spatial and temporal dependencies, significantly boosting performance in human activity recognition. At this stage, it is possible to select the number of antennas used in the training process. The more antennas that are selected—up to the limit of four—the more distinctive information becomes available, consequently improving the neural network’s recognition performance.
To understand how the IBIS operates, it is important to recognize the characteristics and advantages introduced by this new configuration. Unlike the original Inception network, additional layers are incorporated, providing enhanced data interpretation and improving the model’s ability to extract discriminative features.
@@
Figure 7 illustrates the integrated architecture of these two neural structures. While some layers like Normalization, Reshape, and Dropout are omitted for visualization, they are described in the text. Table 2 provides the input and output dimensions and parameter counts for each block. The shaded and hatched purple areas in the figure highlight the novel components of our proposal. Our approach enhances the standard Inception network, which typically only includes activation and convolution layers for learning local features. We added a MaxPooling layer and Batch Normalization to improve convergence and filter relevant features. Furthermore, a BiLSTM network was incorporated to enable the model to learn temporal patterns and motion dynamics from Doppler shifts.
@@
Inception Network: This module is configured to extract spatial features from the Doppler trace data. A description of each block and its function is as follows:
@@
BiLSTM Network: The second part of the architecture is the BiLSTM network, which was included to capture crucial temporal patterns and dependencies from the Doppler traces. These traces are time-series signals that reflect movement and channel variations, often exhibiting complex patterns that the network must recognize.
@@
  • Dense Layer: A fully connected layer that generates the final output, with the number of units depending on the number of output classes (Table 2).
3.1.3 Post-processing
The original SHARP method used a majority voting mechanism among the four receiving antennas to determine the final activity label. This step smoothed outliers and improved robustness.
Figure 2: Architecture of the Inception-BiLSTM hybrid neural network for Human Activity Recognition using Doppler traces, showing all convolutional and attention layers.
Table 2: Inception-BiLSTM Architecture (Compact Notation)
@@ 3.1.3 Post-processing
3.2 Post-Processing Support Vector Machine
The hybrid Inception-BiLSTM neural network is effective at detecting and classifying both global patterns and temporal sequences. However, for accurate classification of CSI-derived data, this architecture alone is insufficient. To address this, an SVM was integrated into the post-processing stage as an ensemble learning strategy. Unlike common approaches where SVM is the primary classifier, this work uses it to refine the final classification results and improve generalization. The SVM’s ability to handle outliers and its use of decision margins contribute to better model generalization, which is particularly useful for Doppler trace data that is prone to noise and distortions. This allows the SVM to refine the final classification results, especially for samples that might be misclassified by the neural network.
A key advantage of SVM lies in the flexibility of its kernel functions. The Radial Basis Function (RBF) kernel, as well as the Sigmoid and Polynomial kernels [16][17][18], are available options that can be selected for classification. In our framework, during the training stage, the SVM is trained using the probability outputs generated by the neural network. Through a grid-search optimization process, the most suitable kernel is determined for the underlying class distribution.
The penalty parameter C is also tuned to achieve an optimal balance between training accuracy and model generalization. Furthermore, class balancing was applied to the SVM configuration to ensure that all classes receive equal weighting, preventing the classifier from becoming biased toward majority classes and resulting in a fairer and more robust overall prediction performance.
Figure 3 illustrates the SVM model’s application, showing five distinct classes separated by well-defined boundaries. Proper class segmentation is essential for understanding the unique features of each class, leading to improved classification accuracy.
Regarding the hyperparameters, a learning rate of 0.001 was used for the hybrid network, along with an early stopping callback function to achieve the highest possible accuracy. At the classification stage, the SVM was also trained using a Grid Search, where three parameters were varied to determine the optimal training configuration. Initially, three kernel functions were considered: RBF, sigmoid, and polynomial, along with the trade-off parameter $C$ and gamma ($\gamma$), both ranging from 0.01 to 1, which are responsible for defining the decision margins. The RBF kernel was ultimately selected by the Grid Search, representing the non-linear characteristics of the data.
The main reason for incorporating an SVM in the post-processing stage is that the Softmax layer of a neural network tends to force a probabilistic decision, even in highly ambiguous classes, and is sensitive to small variations in previously unseen environments. In contrast, the SVM aims to maximize the separation margin using high-level feature vectors, receiving a cleaner representation of class boundaries. This makes the inclusion of an SVM particularly suitable for noisy Doppler traces, where decision boundaries are not well defined. In this context, the combination of the BiLSTM module with dropout regularization, followed by the SVM in the post-processing step, contributes significantly to improving the model’s generalization capability.